4. Let X,,X, be two r.v.s with m.g.f. given by M(t,,ty)=[} (+2 +1) ++ (e? +e?)]”, tų, tz € R.

CALCULATE C(X1, X2)
Answer #1

solution! Let x, xn be two rivis. with mof MCs, (e) = ( Cette 1) telettet) i Liter F%*) - (691627/5-6. ? ? [$(&#4,1)++ (te!)

E(X) = = 2(6+) (etz) 160 (1+i). E (X₂) = CONCX, XL ] = E(X, Zz] - E(X, JE(X) Z-1(1) = 3-1 = 16 CONCT, X2 ) = 5 /